OpenStudy (anonymous):

find all solutions of sec^2 x + tan^2 x= 3 in the interval [0,2pi)

OpenStudy (anonymous):

sec^2 x -1 = tan^2 x

OpenStudy (mathmate):

sec^2x+tan^2x=3 (1+tan^2x)+tan^2x=3 2tan^2x=3-1 tan^2x=1 tan(x)=+/- 1 Can you take it from here?

OpenStudy (anonymous):

nope its tanx=1 or -1

OpenStudy (anonymous):

Yea, I deleted it. Haha. Sorry, my mistake.

OpenStudy (mathmate):

Sorry that it wasn't clear, tan(x)=+/- 1 was meant to be:\[\tan(x) = \pm 1\]

OpenStudy (anonymous):

45, 135, 225, & 315 degrees, or pi/4, 3pi/4, 5pi/4, & 7pi/4 radians. This is the answer and continuation to mathmates' solution.

OpenStudy (mathmate):

Thank you!